find the x intercept, if any, and y intercept, if any, of the line, use the intercepts to sketch a graph of the line and find the slope of the line.
4x - 3y = -24

Answer by ewatrrr(24785)

 
Hi,
find the x intercept, if any, and y intercept, if any, of the line, use the intercepts
to sketch a graph of the line and find the slope of the line.
4x - 3y = -24 |y-intercept is (0,8) and x-intercept is (-6,0)
Using the standard slope-intercept form for an equation of a line y = mx + b where m is the slope and b the y-intercept.
4x - 3y = -24 OR y = (4/3)x + 8
Note: as y increases by 4, x increases by 3.